An infinite G.P. has its first term as x and sum as 5. Then what is the range of x? 
 

Grade:12th Pass

8 Answers

Arun
6 years ago
First term = x
Sum of infinte terms of a G.P. = 5
Hence
If the common ratio is r, then for a sum to infinity -1
and 5=x/(1-r) so r=1- x/5 giving 
-1
0
ASh
6 years ago
so what is the range?????????????????????????????????????????????????????????????
??????????????????????????????????
Arun
6 years ago
If the common ratio is r, then for a sum to infinity -1
and 5=x/(1-r) so r=1- x/5 giving 
-1
0
Arun
6 years ago
If the common ratio is r, then for a sum to infinity -1
and 5=x/(1-r) so r=1- x/5 giving 
-1
 
0
 
I dont know why my comment is not posted fully
ASh
6 years ago
i can see only -1 and 0
that is
here is what i see
 
If the common ratio is r, then for a sum to infinity -1
and 5=x/(1-r) so r=1- x/5 giving 
-1
 
0
 
I dont know why my comment is not posted fully
 
 
ASh
6 years ago
so i am not dissapproving your answer
….........................................................................................................................
Arun
6 years ago
Sum of infinite series
5 = x/1-r
1-r = x/5
r = 1-x/5
-1
Solve it
0
Hope this time you get the full answer.
 
Lucky
6 years ago
Since Sun of infinite GP is x/1-r = 5, where r is a common ratio. Therefore, x=5(1-r) - `a`we know that, -1-r>-1 (multiply by -1)2>1-r>0 (adding 1)10>(1-r)5>0 (multiply by 5 )(from -a)10>x>0 is the required range.